a circuit consists of a 12.0v battery connected to three resistors (42 ohms, 17 ohms, and 110 ohms) in series. Find (a) the current that flows through the battery and (b) the potential difference across each resistor.
Answer #1

the combined resistance is

R = R1 + R2 + R3 = 42+17+110 = 169 ohms

Current in the circuit is

I = V/ R = 12 /169 = 0.071 A

b) Potential difference across the resistor is

V_1 = IR_1 = (0.071)(42) = 2.982 volts

V_2 = IR_2= (0.071)(17) = 1.207 volts

V_3 = IR_3= (0.071)(110) = 7.81 volts